# Optimal Convergence Rate for Mirror Descent Methods with special Time-Varying Step Sizes Rules

### v3

@@ 6. Numerical experiments
To show the advantages of the proposed algorithms and effects of the weighting scheme for generated points $\hat{x}$ in Theorem 3.1 (and the same points in Theorem 4.3), in the convergence of the studied algorithms, a series of numerical experiments were performed for some problems with a geometrical nature in addition to the problem of minimizing the maximum of a finite collection of linear functions.
We compare the performance of the studied algorithms with the projected subgradient method using different famous step size rules that are listed in Table 1 (which is inspired from [16]). Therefore, in our experiments, we take the standard Euclidean prox-structure, namely $\psi(x)=\frac{1}{2}\|x\|_{2}^{2}$ which is $1$-strongly functions (i.e., $\sigma=1$) and the corresponding Bergman divergence is $V_{\psi}(x,y)=\frac{1}{2}\|x-y\|_{2}^{2}$. In all experiments, we take the set $Q$ as a unit ball in $\mathbb{R}^{n}$ with the center at $0\in\mathbb{R}^{n}$. All compared methods start from the same initial point $x^{1}=\left(\frac{1}{\sqrt{n}},\ldots,\frac{1}{\sqrt{n}}\right)\in Q\subset\mathbb{R}^{n}$. In the AdaGrad algorithm, there is an assumption that $\|x^{1}-x^{*}\|_{2}^{2}\leq 2\theta_{0}^{2}$, thus for the taken feasible set in our experiments, we can take $\theta_{0}=\sqrt{2}$.
The comparison of the methods is done in terms of the difference $\hat{f}_{k}-f_{\text{min}}$, where $\hat{f}_{k}$ denotes the value of the objective function $f$ at the averaged points (namely at $\hat{x}_{k}=\frac{1}{k}\sum_{j=1}^{k}x^{k}$, for all cases in Table 1, except to the case of ”quad grad”, where in this case we have $\hat{x}_{k}=\left(\sum_{j=1}^{k}\gamma_{k}\right)^{-1}\sum_{j=1}^{k}\gamma_{k}x^{k}$) and $f_{\text{min}}$ denotes the minimal value of the objective function computed by SciPy, a package for solving many different classes of optimization problems (when the dimension of the space $\mathbb{R}^{n}$ is not big (for example when $n\leq 200$)). Also, we did a comparison of the methods when $\hat{f}_{k}$ denotes the best value of $f(x^{k})$ attained, i.e., $\hat{f}_{k}=\min\{f(x^{1}),\ldots,f(x^{k})\}$, see Remarks 2 and 5.
@@
constant step constant step size [11] $\gamma_{k}=0.1$
fixed length fixed step length [11] $\gamma_{k}=\frac{0.2}{\|\nabla f(x^{k})\|_{*}}$
nonsum non-summable diminishing step [11] $\gamma_{k}=\frac{0.1}{\sqrt{k}}$
sqrsum nonsum square summable but not summable step [11] $\gamma_{k}=\frac{0.5}{k}$
quad grad quadratic of the norm of the gradient [36] $\gamma_{k}=\frac{0.2}{\|\nabla f(x^{k})\|_{*}^{2}}$
AdaGrad AdaGrad algorithm [15] $\gamma_{k}=\frac{\theta_{0}}{\sqrt{\sqrt{\sum_{j=1}^{k}\|\nabla f(x^{j})\|_{*}^{2}+\alpha}}}$
Polyak step Polyak step size [29] $\gamma_{k}=\frac{f(x^{k})-f^{*}}{\|\nabla f(x^{k})\|_{*}^{2}}$
new non-adaptive this paper (3.7) $\gamma_{k}=\frac{\sqrt{2\sigma}}{M_{f}\sqrt{k}}$
new adaptive this paper (3.7) $\gamma_{k}=\frac{\sqrt{2\sigma}}{\|\nabla f(x^{k})\|_{*}\sqrt{k}}$
@@ 6.1. Best approximation problem
The point $A$ is randomly generated from a uniform distribution over $[0,1)$, such that $\|A\|_{2}=10$, therefore the distance between the point $A$ and the considered unit ball $Q$ is equal to $9$, i.e., $f^{*}=9$. Here, we mention that this problem is constructed so that it can be used the Polyak step size, which requires knowing the optimal value $f^{*}$. The results of the comparison, for problem (6.1) with $n=1000$ are presented in Fig. 1. In this figure $\hat{f}^{k}$ denotes the value of the objective function $f$ at the averaged points in each iteration of all compared algorithms. From this figure, we can see that Algorithm 1, with non-adaptive and adaptive step size rules (3.7), works better than other algorithms, where the difference between the performance of Algorithm 1 and the rest of algorithms with steps in Table 1 is clear and significant. We also note that there is no significant difference in the performance of the Algorithm 1 when using the adaptive or non-adaptive step size rules steps
Figure 1. Results of Algorithm 1 for problem (6.1) with $n=1000$.
@@
Figure 2. Results of Algorithm 1 for problem (6.1) with $n=1000$. Here $\hat{f}_{k}=\min\{f(x^{1}),\ldots,f(x^{k})\}$.
@@
Figure 3. Results of Algorithm 1 for problem (6.1) with $n=1000$ with different values of the parameter $m$.
@@
The results of the comparison, for problem (6.2) with $n=200$ and $T=25$, are presented in Fig. 4. The reason here for taking $n=200,T=25$ is that we calculated the value $f_{\text{min}}$ using SciPy which does not well work for large values. In Fig. 4, $\hat{f}^{k}$ denotes the value of the objective function $f$ at the averaged points in each iteration of all compared algorithms. From this figure, we can see that Algorithm 1, with non-adaptive and adaptive step size rules (3.7), outperforms the other methods providing a better solution.
Figure 4. Results of Algorithm 1 for problem (6.2) with $n=200,T=25$.
@@
Figure 5. Results of Algorithm 1 for problem (6.2) with $n=200,T=25$. $\hat{f}_{k}=\min\{f(x^{1}),\ldots,f(x^{k})\}$.
@@
Figure 7. Results of Algorithm 1 for problem (6.2) with $n=1000,T=100$.
@@ 6.3. Smallest covering ball problem
Figure 8. Results of Algorithm 1 for problem (6.3) with $n=200,T=25$ (in the left). And with $n=1000,T=100$ (in the right).
@@ 6.4. Maximum of a finite collection of linear functions
Figure 9. Results of Algorithm 1 for problem (6.4) with $n=200,T=25$ (in the left). And with $n=1000,T=100$ (in the right).
@@ 6.5. Problems with functional constraints
Figure 10. Results of Algorithm 3 and Algorithms 1, 2 in [3], for problem (6.1) (in the left) and problem (6.2) (in the right), with functional constraints (6.5), and $n=1000,p=100,T=100$.
Figure 11. Results of Algorithm 3 and Algorithms 1, 2 in [3], for problem (6.3) (in the left) and problem (6.4) (in the right), with functional constraints (6.5), and $n=1000,p=100,T=100$.

### v5

@@ 6. Numerical experiments
To show the advantages and effects of the weighting scheme for generated points $\hat{x}$, by Algorithm 1 (see Theorem 3.1) (the same will be for the generated points $\hat{x}$ by Algorithm 2 (see Theorem 4.3)), in the convergence of algorithms, a series of numerical experiments were performed for some problems with a geometrical nature in addition to the problem of minimizing the maximum of a finite collection of linear functions.
We compare the performance of the studied algorithms with the projected subgradient method using different famous step size rules that are listed in Table 1. Therefore, in our experiments, we take the standard Euclidean prox-structure, namely $\psi(x)=\frac{1}{2}\|x\|_{2}^{2}$ which is $1$-strongly functions (i.e., $\sigma=1$) and the corresponding Bregman divergence is $V_{\psi}(x,y)=\frac{1}{2}\|x-y\|_{2}^{2}$. In all experiments, we take the set $Q$ as a unit ball in $\mathbb{R}^{n}$ with the center at $0\in\mathbb{R}^{n}$. All compared methods start from the same initial point $x^{1}=\left(\frac{1}{\sqrt{n}},\ldots,\frac{1}{\sqrt{n}}\right)\in Q\subset\mathbb{R}^{n}$. In the AdaGrad algorithm, we take $\alpha=10^{-8}$, and there is an assumption that $\|x^{1}-x^{*}\|_{2}^{2}\leq 2\theta_{0}^{2}$, thus for the taken feasible set in our experiments, we can take $\theta_{0}=1/\sqrt{2}$.
The comparison of the methods is done in terms of the difference $\hat{f}_{k}-f_{\text{min}}$, where $\hat{f}_{k}$ denotes the value of the objective function $f$ at the averaged points (namely at $\hat{x}_{k}=\frac{1}{k}\sum_{j=1}^{k}x^{k}$, for all cases in Table 1, except to the case of ”quad grad”, where in this case we have $\hat{x}_{k}=\left(\sum_{j=1}^{k}\gamma_{k}\right)^{-1}\sum_{j=1}^{k}\gamma_{k}x^{k}$) and $f_{\text{min}}$ denotes the minimal value of the objective function computed by SciPy, a package for solving many different classes of optimization problems (when the dimension of the space $\mathbb{R}^{n}$ is not big. Also, we did a comparison of the methods when $\hat{f}_{k}$ denotes the best value of $f(x^{k})$ attained, i.e., $\hat{f}_{k}=\min\{f(x^{1}),\ldots,f(x^{k})\}$, see Remarks 2 and 5.
@@
constant step constant step size [11] $\gamma_{k}=0.1$,
fixed length fixed step length [11] $\gamma_{k}=\frac{0.2}{\|\nabla f(x^{k})\|_{*}}$,
nonsum non-summable diminishing step [11] $\gamma_{k}=\frac{0.1}{\sqrt{k}}$,
sqrsum nonsum square summable but not summable step [11] $\gamma_{k}=\frac{0.5}{k}$,
quad grad quadratic of the norm of the gradient [35] $\gamma_{k}=\frac{0.2}{\|\nabla f(x^{k})\|_{*}^{2}}$,
AdaGrad AdaGrad algorithm [15] $\gamma_{k}=\frac{\theta_{0}}{\sqrt{\sum_{j=1}^{k}\|\nabla f(x^{j})\|_{*}^{2}+\alpha}}$,
Polyak step Polyak step size [28] $\gamma_{k}=\frac{f(x^{k})-f^{*}}{\|\nabla f(x^{k})\|_{*}^{2}}$,
new non-adaptive this paper (3.7) $\gamma_{k}=\frac{\sqrt{2\sigma}}{M_{f}\sqrt{k}}$,
new adaptive this paper (3.7) $\gamma_{k}=\frac{\sqrt{2\sigma}}{\|\nabla f(x^{k})\|_{*}\sqrt{k}}$.
@@ 6.1. Best approximation problem
The point $A$ is randomly generated from a uniform distribution over $[0,1)$, such that $\|A\|_{2}=10$, therefore the distance between the point $A$ and the considered unit ball $Q$ is equal to $9$, i.e., $f^{*}=9$. Here, we mention that this problem is constructed so that it can be used the Polyak step size, which requires knowing the optimal value $f^{*}$. The results of the comparison, for problem (6.1) with $n=1000$ are presented in Fig. 1. In this figure $\hat{f}_{k}$ denotes the value of the objective function $f$ at the averaged points in each iteration of all compared algorithms. From this figure, we can see that Algorithm 1, with non-adaptive and adaptive step size rules (3.7), works better than other algorithms, where the difference between the performance of Algorithm 1 and the rest of algorithms with steps in Table 1 is clear and significant. We also note that there is no significant difference in the performance of the Algorithm 1 when using the adaptive or non-adaptive step size rules steps
Figure 1. Results of Algorithm 1 and projected subgradient method using different step size rules listed in Table 1, for problem (6.1) with $n=1000,m=5$.
@@
Figure 2. Results of Algorithm 1 and projected subgradient method using different step size rules listed in Table 1, for problem (6.1) with $n=1000,m=5$. Here $\hat{f}_{k}=\min\{f(x^{1}),\ldots,f(x^{k})\}$.
@@
Figure 3. Results of Algorithm 1 for problem (6.1) with $n=1000$ and different values of the parameter $m$.
@@
The results of the comparison, for problem (6.2) with $n=200$ and $T=25$, are presented in Fig. 4. The reason here for taking $n=200,T=25$ is that we calculated the value $f_{\text{min}}$ using SciPy which does not work well for large values. In Fig. 4, $\hat{f}_{k}$ denotes the value of the objective function $f$ at the averaged points in each iteration of all compared algorithms. From this figure, we can see that Algorithm 1, with non-adaptive and adaptive step size rules (3.7), outperforms the other methods providing a better solution.
Figure 4. Results of Algorithm 1 and projected subgradient method using different step size rules listed in Table 1, for problem (6.2) with $n=200,T=25,m=5$.
@@
Figure 5. Results of Algorithm 1 and projected subgradient method using different step size rules listed in Table 1, for problem (6.2) with $n=200,T=25,m=5$. Here $\hat{f}_{k}=\min\{f(x^{1}),\ldots,f(x^{k})\}$.
@@
Figure 7. Results of Algorithm 1 and projected subgradient method using different step size rules listed in Table 1, for problem (6.2) with $n=1000,T=100,m=5$.
@@ 6.3. Smallest covering ball problem
Figure 8. Results of Algorithm 1 and projected subgradient method using different step size rules listed in Table 1, for problem (6.3) with $n=200,T=25,m=5$ (left). And with $n=1000,T=100,m=5$ (right).
@@ 6.4. Maximum of a finite collection of linear functions
Figure 9. Results of Algorithm 1 and projected subgradient method using different step size rules listed in Table 1, for problem (6.4) with $n=200,T=25,m=5$ (left). And with $n=1000,T=100,m=5$ (right).
@@ 6.5. Problems with functional constraints
Figure 10. Results of Algorithm 3 and Algorithms 1, 2 in [3], for problem (6.1) (left) and problem (6.2) (right), with functional constraints (6.5), and $n=1000,p=100,T=100$.
Figure 11. Results of Algorithm 3 and Algorithms 1, 2 in [3], for problem (6.3) (left) and problem (6.4) (right), with functional constraints (6.5), and $n=1000,p=100,T=100$.